# The stable trees revisited

### v1

@@ 3.3 Relation to size-biased distributions
In light of the previous section, we seek to understand the law of the size-biased random reordering $\hat{\bm{D}}$. We will show that this vector is related by a measure change to an i.i.d. sequence $\xi^{*}$ of random variables with the size-biased distribution: that is, $\xi_{1}^{*},\xi_{2}^{*},\dots,\xi_{n}^{*}$ are i.i.d. with law
@@
We will mostly be interested in the first $m$ entries of these vectors, where $m=o(n)$ as $n\to\infty$. We can only relate the laws on the event $\{N^{n}\geq m\}$, where $N^{n}$ is the number of non-zero entries of $\hat{\bm{D}}$, but in our case this event has probability tending to 1. In what follows, write $\Xi_{n}$ to denote a sum of $n$ i.i.d. copies of $\xi$ (note that this is a sum of non-size-biased offspring distributions).
@@
For $1\leq m<n$ and any bounded continuous test function $f:\mathbb{R}^{m}\to\mathbb{R}$, we have
@@
By linearity it is enough to consider the case $f(x_{1},\dots,x_{m})=\mathbbm{1}\left(x_{1}=r_{1},\dots,x_{m}=r_{m}\right)$, where $r_{1},\dots,r_{m}$ are all at least 1 and sum to at most n1n-1. For such a choice of $f$ we have
@@
As an example, consider the length $Y_{1}^{n}$ of the first stick in the construction. (We define the length of a stick to mean the number of edges it contains). We note that $Y_{1}^{n}=C_{1}-1$, so in particular the first stick has length at least $k$ (i.e. it consists of $k+1$ vertices and $k$ edges) iff the first $k$ steps in the algorithm are all growth events. Conditionally given $\hat{\bm{D}}$, we see
@@ 3.4 Heuristics for scaling limits in the $\alpha$-stable case
and now, by the local limit theorem, if $\frac{1}{n^{1/\alpha}}\sum_{i=1}^{m}(\xi_{i}^{*}-1)$ converges in distribution to some random variable $Z$, then the above probability is asymptotic to $n^{-1/\alpha}p(-Z)$. Meanwhile the probability $\mathbb{P}\left(\Xi_{n}=n-1\right)$ is asymptotic to $n^{-1/\alpha}p(0)$, and so the ratio converges to p(Z)/p(0)p(-Z)/p(0).
@@
Assuming that branching events are rare in the discrete process, the partial sums of $\hat{\bm{D}}$ approximately give the number of free half-edges attached to the subtree already created at each step of the construction. Near the beginning of the process, the probability of a branching event is roughly $1/n$ times this number of free half-edges, and so $\hat{\bm{D}}$ is playing an analogous role to the intensity function in a line-breaking construction. Thus, $\widetilde{\sigma}$ acts as the intensity for a line-breaking construction that yields the $\alpha$-stable tree. Indeed, looking back to our first stick length example, if we take $k=\lfloor sn^{1-\frac{1}{\alpha}}\rfloor$ for some $s\leq t$, then we see
@@
and by similar calculations to the ones used for $\Theta_{m}^{n}$, the product term should converge to $\exp(-\int_{0}^{s}\sigma_{s}\,ds)$. As the $\Theta$ term converges to the measure change, the limit of this expectation should be $\mathbb{E}\left[\exp(-\int_{0}^{s}\widetilde{\sigma}_{s}\,ds)\right]$, which is exactly the probability that the first stick has length $\geq s$ in the line-breaking construction driven by $\widetilde{\sigma}$.
@@
Recall that we assume the offspring distribution $\xi$ is in the domain of attraction of a stable law of index $\alpha$. As in the previous section, we let $\Xi_{n}=\sum_{i=1}^{n}\xi_{i}$, where $\xi_{i}$ are i.i.d. copies of $\xi$. By Theorem 1 of Kortchemski [igorduquesneproof], there exists an increasing sequence $a_{n}>0$, regularly varying of index $1/\alpha$, such that $\frac{\Xi_{n}-n}{a_{n}}$ converges in distribution to $L_{1}$, where $L$ is the spectrally positive $\alpha$-stable Lévy process, and moreover we have the local limit theorem
@@
As $\xi$ is in the domain of attraction of an $\alpha$-stable law, there exists a slowly varying function $l(t)$ such that $\mathbb{P}\left(\xi\geq t\right)=l(t)t^{-\alpha}$ for all $t>0$. By the proof used in Theorem 3.8.2 in [durrettprobth], we also know $n\mathbb{P}\left(\xi\geq xa_{n}\right)\to\int_{x}^{\infty}Ct^{-\alpha-1}\,dt=\frac{C}{\alpha}x^{-\alpha}$ for any fixed $x$. Combining these results, we deduce that
$l(xa_{n})\sim\frac{Ca_{n}^{\alpha}}{\alpha n}$
@@
and so by Karamata’s integral theorem (Proposition 1.5.10 of [RegularVariation]) we have $\mathbb{P}\left(\xi^{*}\geq t\right)\sim\frac{\alpha}{\alpha-1}l(t)t^{-\alpha+1}$. In particular, for fixed $x$ we have
$\mathbb{P}\left(\xi^{*}\geq xa_{n}\right)\sim\frac{\alpha}{\alpha-1}l(xa_{n})(xa_{n})^{-\alpha+1}\sim\frac{1}{m_{n}}\frac{C}{\alpha-1}x^{-\alpha+1}=\frac{1}{m_{n}}\int_{x}^{\infty}t^{-\alpha}\,dt$
and so, following the proof in Theorem 3.8.2 in [durrettprobth], we deduce the desired convergence in distribution. ∎
@@ 4.1 Convergence of the measure change
$\prod_{i=1}^{\lfloor tm_{n}\rfloor}\left(\frac{n-i+1}{n-1-\sum_{j=1}^{i-1}\xi_{i}^{n,*}}\right)$ $=\prod_{i=1}^{\lfloor tm_{n}\rfloor}\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)-1}{n-i+1}\right)^{-1}$
$=\exp\left(-\sum_{i=1}^{\lfloor tm_{n}\rfloor}\log\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)-1}{n-i+1}\right)\right)$
$=\exp\left(-\int_{0}^{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}m_{n}\log\left(1-\frac{\sum_{j=1}^{\lfloor sm_{n}\rfloor}(\xi_{j}^{n,*}-1)-1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right)$
$=\exp\left(-\int_{0}^{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)-1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right).$
@@
$\left(m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)-1}{n-\lfloor sm_{n}\rfloor}\right)+\sigma^{n}(s)\right)_{s\in[0,T]}\to 0\quad\text{ a.s.\ in the Skorokhod metric,}$
@@
$\int_{0}^{t}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)-1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\to\int_{0}^{t}\sigma_{s}\,ds$
@@
$\sup_{t\in[0,T]}\left|\int_{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}^{t}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)-1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right|\to 0\quad\text{ a.s.}$
@@
$\left|\frac{a_{n}}{a_{n-k}}p\left(-\frac{a_{n}}{a_{n-k}}\sigma^{n}(t)-\frac{1}{a_{n-k}}\right)-p\left(-\frac{a_{n}}{a_{n-k}}\sigma^{n}(t)-\frac{1}{a_{n-k}}\right)\right|\leq\sup_{t\in[0,T]}\left\{\frac{a_{n}}{a_{n-k(t)}}-1\right\}\|p\|_{\infty}.$
@@
Note that $\mathbb{E}\left[|M_{t}^{n}|\right]=\mathbb{E}\left[M_{t}^{n}\right]=\mathbb{P}\left(N^{n}\geq\lfloor Tn^{1-\frac{1}{\alpha}}\rfloor\right)\to 1=\mathbb{E}\left[|M_{t}|\right]$, and so by Scheffé’s lemma we have $\mathbb{E}\left[|M_{n}^{t}-M_{t}|\right]\to 0$ in the setting that everything is defined on the same space.
@@ 4.2 Convergence of random finite-dimensional distributions
where $Y_{i}$ are the jump times of a Poisson process of intensity $\widetilde{\sigma}$, and conditionally given $\widetilde{\sigma}$ and $(Y_{1},\dots,Y_{k})$, the random variables $(Z_{1},\ldots,Z_{k})$ are independent and $Z_{i}$ has c.d.f. $\frac{\widetilde{\sigma}_{t}}{\widetilde{\sigma}_{Y_{i}-}},t\in[0,Y_{i})$. Our strategy of proof is similar to that of Section 4 of [snakes].
@@
Let $k\geq 0$ and let $t_{1},\dots,t_{k}\leq T$ be fixed. Define processes $Q^{n,k}_{t},Q^{k}_{t}$ for $t\in[0,T]$ by
@@
$Q_{t}^{k}=\sigma_{t-}\exp\left(-\int_{0}^{t}\sigma_{u}\,du\right)\int_{\prod_{i=1}^{k}[0,t_{i}]}\mathbbm{1}\left(u_{1}<u_{2}<\dots<u_{k}<t\right)\prod_{i=1}^{k}\left(\sigma_{u_{i}^{-}}\wedge\sigma_{s_{i}}\right)\,du_{1}\dots du_{k},$
@@
Now adding in attachment points $\zeta_{1},\zeta_{2},\dots$, again such that conditionally given $\sigma$ and $(\eta_{j})_{j\geq 1}$, the random variables $(\zeta_{1},\zeta_{2},\ldots)$ are independent and $\zeta_{i}$ has c.d.f. $\sigma_{t}/\sigma_{\eta_{i}^{-}},t\in[0,\eta_{i})$, we obtain
$\mathbb{P}\left(\eta_{1}\in du_{1},\dots,\eta_{k}\in du_{k},\zeta_{1}\leq s_{1},\dots,\zeta_{k-1}\leq s_{k-1}\,\big|\,\sigma\right)\\ =\mathbbm{1}\left(u_{1}<\dots<u_{k}\right)\sigma_{u_{k}}\prod_{i=1}^{k-1}\left(\left(\sigma_{s_{i}}\wedge\sigma_{u_{i}^{-}}\right)\frac{\sigma_{u_{i}}}{\sigma_{u_{i}^{-}}}\right)\exp\left(-\int_{0}^{u_{k}}\sigma_{r}\,dr\right)\,du_{1}\dots du_{k}.$
@@
$\mathbb{P}\left(\eta_{1}\leq t_{1},\dots,\eta_{k}\leq t_{k},\eta_{k+1}\in dt,\zeta_{1}\leq s_{1},\dots,\zeta_{k}\leq s_{k}\right)\\ =dt\,\sigma_{t}\exp\left(-\int_{0}^{t}\sigma_{u}\,du\right)\int_{\prod_{i=1}^{k}[0,t_{i}]}\mathbbm{1}\left(u_{1}<\dots<u_{k}<t\right)\prod_{i=1}^{k}\left(\left(\sigma_{s_{i}}\wedge\sigma_{u_{i}^{-}}\right)\frac{\sigma_{u_{i}}}{\sigma_{u_{i}^{-}}}\right)\,du_{1}\dots du_{k}$
which is almost in the form of $Q_{t}^{k}\,dt$, except for the factors of $\sigma_{u_{i}}/\sigma_{u_{i}^{-}}$ in the integral. However, $\sigma$ has countably many discontinuities, and so $\prod_{i=1}^{k}\frac{\sigma_{u_{i}}}{\sigma_{u_{i}^{-}}}=1$ Lebesgue-almost everywhere on $\prod_{i=1}^{k}[0,t_{i}]$. Thus we have exactly
@@
We will use the Skorokhod representation again and prove almost sure convergence in the Skorokhod metric. We introduce auxiliary processes $R_{t}^{n,k}$, defined analogously to $Q_{t}^{k}$ but with $\sigma$ replaced by $\sigma^{n}$. With this we know $R^{n,k}\to Q^{k}$ in the Skorokhod metric (using the same time change that gives $\sigma^{n}\to\sigma$), and so it is enough to show that as $n\to\infty$, the difference between $Q_{t}^{n,k}$ and $R_{t}^{n,k}$ tends to 0 uniformly in $t$. We seek to prove this stronger convergence result by induction on $k$.
@@
$\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{F}_{k}^{n}\right)=\frac{\sum_{j=1}^{r-1-B(r)}(\xi^{n,*}_{i}-1)-B(r)}{\sum_{j=1}^{C^{n}_{k}-k}(\xi_{i}^{n,*}-1)-(k-1)},\quad 0\leq r\leq C^{n}_{k}-1$
@@
$\frac{\sum_{j=1}^{r-k}(\xi^{n,*}_{j}-1)-(k-1)}{\sum_{j=1}^{C^{n}_{k}-k}(\xi_{j}^{n,*}-1)-(k-1)}\leq\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{F}_{k}^{n}\right)\leq\frac{\sum_{j=1}^{r-1}(\xi^{n,*}_{j}-1)}{\sum_{j=1}^{C_{k}-k}(\xi_{j}^{n,*}-1)-(k-1)}$
@@
$\frac{\sum_{j=1}^{r-k}(\xi^{n,*}_{j}-1)-(k-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}\leq\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{A}\right)\leq\frac{\sum_{j=1}^{r-1}(\xi^{n,*}_{j}-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}.$
@@
which depends on $\mathcal{F}_{k}^{n}$ only through $C_{k}^{n}$. Thus for any $\mathcal{F}_{k}^{n}$-measurable event $\mathcal{A}\subset\{C_{k}^{n}=l\}$ we have
@@
$\widetilde{Q}_{t}^{n,k}$ $=m_{n}\mathbb{P}\left(\mathcal{A}_{k-1},J_{k}\leq\lfloor s_{k}m_{n}\rfloor,C^{n}_{k}\leq\lfloor t_{k}m_{n}\rfloor,C^{n}_{k+1}=\lfloor tm_{n}\rfloor,\mathcal{E}^{n,k}\ \Big|\ \hat{\bm{D}}=\xi^{*}\right)$
@@
$\leq\sum_{l=0}^{\lfloor t_{k}m_{n}\rfloor}\frac{\widetilde{Q}_{l/m_{n}}^{n,k-1}}{m_{n}}\left(\frac{\sum_{j=1}^{\lfloor s_{k}m_{n}\rfloor-1}(\xi_{j}^{n,*}-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}\wedge 1\right)$
$\quad\quad\quad\quad\times\prod_{i=l+1}^{\lfloor tm_{n}\rfloor-1}\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)}{n-i+k}\right)\times\frac{m_{n}\sum_{j=1}^{\lfloor tm_{n}\rfloor-1}(\xi_{j}^{n,*}-1)-m_{n}k}{n-\lfloor tm_{n}\rfloor}.$

### v2

@@ 3.3 Relation to size-biased distributions
In light of the previous section, we seek to understand the law of the size-biased random reordering $\hat{\bm{D}}$. We will show that this vector is related by a measure change to an i.i.d. sequence of random variables $(\xi_{i}^{*})$ with the size-biased distribution: that is, $\xi_{1}^{*},\xi_{2}^{*},\dots,\xi_{n}^{*}$ are i.i.d. with law
@@
We will mostly be interested in the first $m$ entries of these vectors, where $m=o(n)$ as $n\to\infty$. We can only relate the laws on the event $\{N^{n}\geq m\}$, where $N^{n}$ is the number of non-zero entries of $\hat{\bm{D}}$, but in our case this event will turn out to have probability tending to 1. In what follows, write $\Xi_{n}$ to denote a sum of $n$ i.i.d. copies of $\xi$ (note that this is a sum of non-size-biased offspring distributions).
@@
For $1\leq m<n$ and any bounded test function $f:\mathbb{Z}^{m}\to\mathbb{R}$, we have
@@
By linearity it is enough to consider the case
$f(x_{1},\dots,x_{m})=\mathbbm{1}\left(x_{1}=r_{1},\dots,x_{m}=r_{m}\right),$
where $r_{1},\dots,r_{m}$ are all at least 1 and sum to at most n1n-1. For such a choice of $f$ we have
@@
As an example, consider the length $Y_{1}^{n}$ of the first stick in the construction. (We define the length of a stick to mean the number of edges it contains). We note that $Y_{1}^{n}=C_{1}-1$, so in particular the first stick has length at least $k$ iff the first $k$ steps in the algorithm are all growth events. Conditionally given $\hat{\bm{D}}$, we see
@@ 3.4 Heuristics for scaling limits in the $\alpha$-stable case
and now, by the local limit theorem for $\Xi$, if $\frac{1}{n^{1/\alpha}}\sum_{i=1}^{m}(\xi_{i}^{*}-1)$ converges in distribution to some random variable $Z$, then the above probability is asymptotic to $n^{-1/\alpha}p(-Z)$. Meanwhile the probability $\mathbb{P}\left(\Xi_{n}=n-1\right)$ is asymptotic to $n^{-1/\alpha}p(0)$, and so the ratio converges to p(Z)/p(0)p(-Z)/p(0).
@@
Assuming that branching events are rare in the discrete process, the partial sums of $\hat{\bm{D}}$ approximately give the number of free half-edges attached to the subtree present at each step of the construction. Near the beginning of the process, the probability of a branching event is roughly $1/n$ times this number of free half-edges, and so $\hat{\bm{D}}$ is playing an analogous role to the intensity function in a line-breaking construction. Thus, we expect $\widetilde{\sigma}$ to act as the intensity for a line-breaking construction that yields the $\alpha$-stable tree. Indeed, looking back to our first stick length example, if we take $k=\lfloor sn^{1-\frac{1}{\alpha}}\rfloor$ for some $s\leq t$, then we see
@@
and by similar calculations to the ones used for $\Theta_{m}^{n}$, the product term should converge to $\exp(-\int_{0}^{s}\sigma_{r}\,dr)$. As the $\Theta$ term converges to the measure change, the limit of this expectation should be $\mathbb{E}\left[\exp(-\int_{0}^{s}\widetilde{\sigma}_{r}\,dr)\right]$, which is exactly the probability that the first stick has length $\geq s$ in the line-breaking construction driven by $\widetilde{\sigma}$.
@@
Recall that we assume the offspring distribution $\xi$ is in the domain of attraction of a stable law of index $\alpha$. As in the previous section, we let $\Xi_{n}=\sum_{i=1}^{n}\xi_{i}$, where $\xi_{i}$ are i.i.d. copies of $\xi$. By Theorem 1 of Kortchemski [33], there exists an increasing sequence $a_{n}>0$, regularly varying of index $1/\alpha$, such that $\frac{\Xi_{n}-n}{a_{n}}$ converges in distribution to $L_{1}$, where $L$ is the spectrally positive $\alpha$-stable Lévy process, and moreover we have the local limit theorem
@@
As $\xi$ is in the domain of attraction of an $\alpha$-stable law, there exists a slowly varying function $l(t)$ such that $\mathbb{P}\left(\xi\geq t\right)=l(t)t^{-\alpha}$ for all $t>0$. By the proof used in Theorem 3.8.2 in [25], we also know $n\mathbb{P}\left(\xi\geq xa_{n}\right)\to\int_{x}^{\infty}C_{\alpha}t^{-\alpha-1}\,dt=\frac{C_{\alpha}}{\alpha}x^{-\alpha}$ for any fixed $x$. Combining these results, we deduce that
$l(xa_{n})\sim\frac{C_{\alpha}a_{n}^{\alpha}}{\alpha n}$
@@
and so by Karamata’s integral theorem (Proposition 1.5.10 of [15]) we have $\mathbb{P}\left(\xi^{*}\geq t\right)\sim\frac{\alpha}{\alpha-1}l(t)t^{-\alpha+1}$. In particular, for fixed $x$ we have
$\mathbb{P}\left(\xi^{*}\geq xa_{n}\right)\sim\frac{\alpha}{\alpha-1}l(xa_{n})(xa_{n})^{-\alpha+1}\sim\frac{1}{m_{n}}\frac{C_{\alpha}}{\alpha-1}x^{-\alpha+1}=\frac{1}{m_{n}}\int_{x}^{\infty}C_{\alpha}t^{-\alpha}\,dt$
and so, following the proof in Theorem 3.8.2 in [25], we deduce the desired convergence in distribution. ∎
@@ 4.1 Convergence of the measure change
$\prod_{i=1}^{\lfloor tm_{n}\rfloor}\left(\frac{n-i+1}{n-1-\sum_{j=1}^{i-1}\xi_{j}^{n,*}}\right)$ $=\prod_{i=1}^{\lfloor tm_{n}\rfloor}\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)+1}{n-i+1}\right)^{-1}$
$=\exp\left(-\sum_{i=1}^{\lfloor tm_{n}\rfloor}\log\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)+1}{n-i+1}\right)\right)$
$=\exp\left(-\int_{0}^{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}m_{n}\log\left(1-\frac{\sum_{j=1}^{\lfloor sm_{n}\rfloor}(\xi_{j}^{n,*}-1)+1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right)$
$=\exp\left(-\int_{0}^{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)+1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right).$
@@
$\left(m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)+1}{n-\lfloor sm_{n}\rfloor}\right)+\sigma^{n}(s)\right)_{s\in[0,T]}\to 0\quad\text{ a.s.\ in the Skorokhod metric,}$
@@
$\int_{0}^{t}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)+1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\to\int_{0}^{t}\sigma_{s}\,ds$
@@
$\sup_{t\in[0,T]}\left|\int_{\frac{\lfloor tm_{n}\rfloor}{m_{n}}}^{t}m_{n}\log\left(1-\frac{a_{n}\sigma^{n}(s)+1}{n-\lfloor sm_{n}\rfloor}\right)\,ds\right|\to 0\quad\text{ a.s.}$
@@
$\left|\frac{a_{n}}{a_{n-k}}p\left(-\frac{a_{n}}{a_{n-k}}\sigma^{n}(t)-\frac{1}{a_{n-k}}\right)-p\left(-\frac{a_{n}}{a_{n-k}}\sigma^{n}(t)-\frac{1}{a_{n-k}}\right)\right|\leq\sup_{t\in[0,T]}\left|\frac{a_{n}}{a_{n-k(t)}}-1\right|\|p\|_{\infty}.$
@@
Note that $\mathbb{E}\left[|M_{t}^{n}|\right]=\mathbb{E}\left[M_{t}^{n}\right]=\mathbb{P}\left(N^{n}\geq\lfloor Tn^{1-\frac{1}{\alpha}}\rfloor\right)\to 1=\mathbb{E}\left[|M_{t}|\right]$, and so by Scheffé’s lemma we have $\mathbb{E}\left[|M_{t}^{n}-M_{t}|\right]\to 0$ in the setting that everything is defined on the same space.
@@ 4.2 Convergence of random finite-dimensional distributions
where $Y_{i}$ are the jump times of a Poisson process of intensity $\widetilde{\sigma}$, and conditionally given $\widetilde{\sigma}$ and $(Y_{1},\dots,Y_{k})$, the random variables $(Z_{1},\ldots,Z_{k})$ are independent and $Z_{i}$ has c.d.f. $\frac{\widetilde{\sigma}_{t}}{\widetilde{\sigma}_{Y_{i}-}},t\in[0,Y_{i})$. Our strategy of proof is similar to that of Section 4 of [2].
@@
Let $k\geq 0$ and let $s_{1},\dots,s_{k},t_{1},\dots,t_{k}$ be fixed with $s_{i}\leq t_{i}\leq T$ for all $i$. Define processes $Q^{n,k}_{t},Q^{k}_{t}$ for $t\in[0,T]$ by
@@
$Q_{t}^{k}=\sigma_{t-}\exp\left(-\int_{0}^{t}\sigma_{u}\,du\right)\!\int_{\prod_{i=1}^{k}[0,t_{i}]}\!\!\!\mathbbm{1}\left(u_{1}<u_{2}<\dots<u_{k}<t\right)\prod_{i=1}^{k}\left(\sigma_{u_{i}-}\wedge\sigma_{s_{i}}\right)du_{1}\dots du_{k},$
@@
Now adding in attachment points $\zeta_{1},\zeta_{2},\dots$, again such that conditionally given $\sigma$ and $(\eta_{j})_{j\geq 1}$, the random variables $(\zeta_{1},\zeta_{2},\ldots)$ are independent and $\zeta_{i}$ has c.d.f. $\sigma_{t}/\sigma_{\eta_{i}-},t\in[0,\eta_{i})$, we obtain
$\mathbb{P}\left(\eta_{1}\in du_{1},\dots,\eta_{k}\in du_{k},\zeta_{1}\leq s_{1},\dots,\zeta_{k-1}\leq s_{k-1}\,\big|\,\sigma\right)\\ =\mathbbm{1}\left(u_{1}<\dots<u_{k}\right)\sigma_{u_{k}}\prod_{i=1}^{k-1}\left(\left(\sigma_{s_{i}}\wedge\sigma_{u_{i}-}\right)\frac{\sigma_{u_{i}}}{\sigma_{u_{i}-}}\right)\exp\left(-\int_{0}^{u_{k}}\sigma_{r}\,dr\right)\,du_{1}\dots du_{k}.$
@@
$\mathbb{P}\left(\eta_{1}\leq t_{1},\dots,\eta_{k}\leq t_{k},\eta_{k+1}\in dt,\zeta_{1}\leq s_{1},\dots,\zeta_{k}\leq s_{k}\right)\\ =dt\,\sigma_{t}\exp\left(-\int_{0}^{t}\sigma_{u}\,du\right)\!\!\int_{\prod_{i=1}^{k}[0,t_{i}]}\!\!\mathbbm{1}\left(u_{1}<\dots<u_{k}<t\right)\prod_{i=1}^{k}\left(\left(\sigma_{s_{i}}\wedge\sigma_{u_{i}-}\right)\frac{\sigma_{u_{i}}}{\sigma_{u_{i}-}}\right)\!du_{1}\dots du_{k}$
which is almost in the form of $Q_{t}^{k}\,dt$, except for the factors of $\sigma_{u_{i}}/\sigma_{u_{i}-}$ in the integral. However, $\sigma$ has countably many discontinuities, and so $\prod_{i=1}^{k}\frac{\sigma_{u_{i}}}{\sigma_{u_{i}-}}=1$ Lebesgue-almost everywhere on $\prod_{i=1}^{k}[0,t_{i}]$. Thus we have exactly
@@
We will use the Skorokhod representation theorem again and prove almost sure convergence in the Skorokhod metric. We introduce auxiliary processes $R_{t}^{n,k}$, defined analogously to $Q_{t}^{k}$ but with $\sigma$ replaced by $\sigma^{n}$. With this we know $R^{n,k}\to Q^{k}$ in the Skorokhod metric (using the same time change that gives $\sigma^{n}\to\sigma$), and so it is enough to show that as $n\to\infty$, the difference between $Q_{t}^{n,k}$ and $R_{t}^{n,k}$ tends to 0 uniformly in $t$. We seek to prove this stronger convergence result by induction on $k$.
@@
$\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{F}_{k}^{n}\right)=\frac{\sum_{j=1}^{r-B(r)}(\xi^{n,*}_{i}-1)-B(r)}{\sum_{j=1}^{C^{n}_{k}-k}(\xi_{i}^{n,*}-1)-(k-1)},\quad 0\leq r\leq C^{n}_{k}-1$
@@
$\frac{\sum_{j=1}^{r-k+1}(\xi^{n,*}_{j}-1)-(k-1)}{\sum_{j=1}^{C^{n}_{k}-k}(\xi_{j}^{n,*}-1)-(k-1)}\leq\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{F}_{k}^{n}\right)\leq\frac{\sum_{j=1}^{r}(\xi^{n,*}_{j}-1)}{\sum_{j=1}^{C_{k}-k}(\xi_{j}^{n,*}-1)-(k-1)}$
@@
$\frac{\sum_{j=1}^{r-k+1}(\xi^{n,*}_{j}-1)-(k-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}\leq\mathbb{P}\left(J_{k}^{n}\leq r\ \Big|\ \mathcal{A}\right)\leq\frac{\sum_{j=1}^{r}(\xi^{n,*}_{j}-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}.$
@@
which depends on $\mathcal{F}_{k}^{n}$ only through $C_{k}^{n}$. Thus for any $\mathcal{F}_{k}^{n}$-measurable event $\mathcal{A}\subset\{C_{k}^{n}=l\}$ with $l\leq r-1$ we have
@@
$\widetilde{Q}_{t}^{n,k}$ $=m_{n}\mathbb{P}\left(\mathcal{A}_{k-1},J_{k}^{n}\leq\lfloor s_{k}m_{n}\rfloor,C^{n}_{k}\leq\lfloor t_{k}m_{n}\rfloor,C^{n}_{k+1}=\lfloor tm_{n}\rfloor,\mathcal{E}^{n,k}\ \Big|\ \hat{\bm{D}}=\xi^{*}\right)$
@@
$\leq\sum_{l=0}^{\lfloor t_{k}m_{n}\rfloor}\frac{\widetilde{Q}_{l/m_{n}}^{n,k-1}}{m_{n}}\left(\frac{\sum_{j=1}^{\lfloor s_{k}m_{n}\rfloor}(\xi_{j}^{n,*}-1)}{\sum_{j=1}^{l-k}(\xi_{j}^{n,*}-1)-(k-1)}\wedge 1\right)$
$\quad\quad\quad\quad\times\prod_{i=l+1}^{\lfloor tm_{n}\rfloor-1}\left(1-\frac{\sum_{j=1}^{i-1}(\xi_{j}^{n,*}-1)}{n-i+k}\right)\times\frac{m_{n}\sum_{j=1}^{\lfloor tm_{n}\rfloor-1-k}(\xi_{j}^{n,*}-1)-m_{n}k}{n-\lfloor tm_{n}\rfloor}.$